### v1

# MedThink: Explaining Medical Visual Question Answering via Multimodal Decision-Making Rationale

Abstract
Medical Visual Question Answering (MedVQA), which offers language responses to image-based medical inquiries, represents a challenging task and significant advancement in healthcare. It assists medical experts to swiftly interpret medical images, thereby enabling faster and more accurate diagnoses. However, the model interpretability and transparency of existing MedVQA solutions are often limited, posing challenges in understanding their decision-making processes. To address this issue, we devise a semi-automated annotation process to streamlining data preparation and build new benchmark MedVQA datasets R-RAD and R-SLAKE. The R-RAD and R-SLAKE datasets provide intermediate medical decision-making rationales generated by multimodal large language models and human annotations for question-answering pairs in existing MedVQA datasets, i.e., VQA-RAD and SLAKE. Moreover, we design a novel framework which finetunes lightweight pretrained generative models by incorporating medical decision-making rationales into the training process. The framework includes three distinct strategies to generate decision outcomes and corresponding rationales, thereby clearly showcasing the medical decision-making process during reasoning. Extensive experiments demonstrate that our method can achieve an accuracy of 83.5% on R-RAD and 86.3% on R-SLAKE, significantly outperforming existing state-of-the-art baselines. Dataset and code will be released.
1 Introduction
The Medical Visual Equal contribution. # Correspondence author.Question Answering (MedVQA) task is designed to take medical images and specialized clinical queries as inputs, and provide accurate answers with texts. Since the inception of the MedVQA challenge in 2018 [?], there has been a significant surge in interest in exploring the capabilities of MedVQA [?]. Effective MedVQA not only holds the potential to enhance patient engagement, thereby alleviating patient stress, but also assists physicians in clinical diagnosis, thus conserving valuable medical resources and reducing the risk of misdiagnosis [?].
The challenges to resolve the MedVQA tasks are two-fold. On one hand, though there exist a wealth of datasets composed of medical images and text annotations  [??], the decision-making process between the question and answer pairs are usually missing, impeding reliable evaluation of model interpretability. While some recent datasets already incorporated images, specialized medical queries, and answer texts  [??], the corresponding reasoning process to reach certain diagnostic decisions remain unclear, resulting in black-box and clinically inapplicable inference [???].
A straightfoward solution is to integrate expert-level reasoning rationales in these datasets to unravel the underlying reasoning processes. However, manual annotation of such rationales is time-consuming and requires in-depth understanding of medical knowledge, while a fast and reliable rationale annotation framework is still missing [??].
On the other hand, models which can resolve MedVQA tasks in a fast, accurate and interpretable manner is of high necessity in real-world applications. Current MedVQA methods often model this problem by retrieval and train MedVQA models with contrastive or classification objectives. For instance, Nguyen et al. [?] employed a combination of unsupervised convolutional denoising autoencoders (CDAE) and the meta-learning method (MAML) to learn domain specific weight initialization of the MedVQA model on external medical datasets. Moreover, Zhang et al. [?] first implemented contrastive learning in the medical domain, presenting ConVIRT, a methodology that utilizes medical text-image contrastive loss for pretraining medical visual representations. Further, Liu et al. [?] proposed CPRD, a two-stage pre-training framework, leveraging representation distillation and contrastive learning to train the visual encoder for MedVQA system on a large corpus of unlabeled radiological images. The recent PubMedCLIP model [?] pioneers the incorporation of the Contrastive Language-Image Pre-Training [?] into MedVQA tasks by conducting pre-training. In contrast, the remarkable performance of large language models (LLMs) across various natural language processing (NLP) tasks has been extended to text question-answering in healthcare [??] recently. Building upon this, multimodal large language models (MLLMs) [??] further accept both text and image inputs to generate responses, presenting a novel approach to tackling MedVQA tasks. However, applying MLLMs directly to MedVQA tasks in real medical scenarios is impractical due to their high operational costs and significant latency.
In this paper, we aim to address the aforementioned challenges by providing new benchmark datasets and novel MedVQA solutions. We design a semi-automated annotation method that leverages the powerful inference capabilities of MLLMs to assist experts during annotation, significantly improving the efficiency. Through our method, we create the R-RAD and R-SLAKE datasets, based on the existing dataset VQA-RAD [?] and SLAKE [?]. The R-RAD and R-SLAKE datasets provide the intermediate reasoning steps required for medical decision-making, including necessary medical background knowledge and descriptions of medical images, which we term Medical Decision-Making Rationales. Moreover, we design a novel framework to finetune the pretrained generative models, specifically selecting the T5-base architecture [?] as our base architecture due to its practicality in real-world applications. With only 220M parameters, the architecture adeptly performs generative tasks, balancing cost-effectiveness and practical value. By incorporating medical decision-making rationales into the training process, our model outputs not only decision outcomes but also corresponding rationales, thereby clearly showcasing the medical decision-making process during reasoning. Based on different inputs for Medical Decision-Making Rationales during training, we further propose three distinct generative modes: “Explanation”, “Reasoning”, and “Two-Stage Reasoning”. Figure LABEL:fig:_case_study illustrates the outcomes of the MedVQA model trained with “Explanation” method.
Extensive experimental results demonstrate that our method achieved an accuracy of 83.5% on the R-RAD dataset and 86.3% on the R-SLAKE dataset. These results represent significant enhancements over the existing state-of-the-art PubMedCLIP [?] model, with improvements of 4.0% and 3.8% respectively. Ablations on different LLMs such as Gemini [?] further demonstrate the effectiveness of dataset and method. Our contributions are as follows:
  • We develop a semi-automated process for annotating MedVQA data with decision-making rationale. To the best of our knowledge, the R-RAD and R-SLAKE datasets represent the first multimodal MedVQA benchmark datasets that encompass rationales for answers.
  • We propose a lightweight framework with three answering strategies, enabling faster and more accurate multimodal MedVQA with enhanced interpretability.
  • We conduct extensive experiments and ablations that demonstrate the usefulness of the R-RAD and R-SLAKE datasets and superiority of our method.
2 Related Work
2.1 MedVQA
VQA represents a cutting-edge, multimodal task at the intersection of computer vision and natural language processing, drawing significant attention in both domains. MedVQA applies the principles of VQA to interpret and respond to complex inquiries about medical imagery. A MedVQA system usually consists of three key components for feature extraction, feature fusion and answer reasoning, respectively, which aims to generate answers in text by processing given medical images. Previous MedVQA solutions [??????] have relied on the CNNs, such as those pretrained on ImageNet like VGGs or ResNets, to extract visual features. Meanwhile, the RNNs are employed to process textual information. With the development of large-scale pretraining, recent works [???] have shifted towards the transformer-based models to enhance feature extraction capabilities for both textual and visual modalities. In terms of content, these works still treat the MedVQA as the classification problem. However, this approach is misaligned with the realities of medical practice, where clinicians rarely face scenarios that can be addressed with predefined answer options.
This incongruity underscores the necessity for a MedVQA approach that is more adaptive and reflective of the complexities inherent in medical diagnostics and decision-making. In this paper, we have redefined MedVQA as the generative task. Within the actual medical environment, when faced with open-ended queries, our proposed MedVQA model can still generate informed responses based on the medical knowledge it has learned.
2.2 The Thought Chain
Recently, NLP has been significantly transformed by language models [???].
To further enhance the reasoning capabilities of language models, prior works [??] have incorporated reasoning rationales during training or inference phases, which guide models to generate the final prediction. On the other hand, in the realm of VQA, it is crucial for VQA systems to understand multimodal information from diverse sources and reason about domain-specific questions. To achieve this goal, several works [??] have proposed multimodal reasoning methods for VQA. These methods, commonly referred to as “the thought chain”, introduces intermediate steps to assist the model in reasoning. In this paper, we present the “Medical Decision-Making Rationale” and apply it to the MedVQA task. We anticipate that MedVQA systems, equipped with the “Medical Decision-Making Rationale”, will not only offer support in medical decision-making but also elucidate the underlying rationales behind these decisions.
3 Methodology
3.1 Dataset Collection
We estalish two benchmark datasets R-RAD and R-SLAKE on top of the VQA-RAD dataset[?] and the SLAKE dataset[?], respectively. VQA-RAD, a meticulously curated medical dataset, sources its radiographic images from MedPix®, an open-access radiology database. In this dataset, clinicians formulate perinent medical questions based on the radiographic images, and provide corresponding answers. The VQA-RAD dataset comprises a collection of 315 images and 3,515 questions, which are categorized as either “closed-end” or “open-end”, depending on the scope of the answer. We adhere to the official dataset split for evaluation.
The SLAKE dataset derives its data from three distinct sources: the ChestX-ray8 [?], the CHAOS Challenge [?], and the Medical Segmentation Decathlon (MSD) [?]. After screening and annotation by medical experts, it yields a bilingual (English-Chinese) MedVQA dataset, including 642 medical images and approximately 14,000 medical questions. These questions are classified as “closed-end” if they have limited answer choices, and “open-end” otherwise. For our work, we utilize only the “English” component of the dataset and follow the original split. After completing the data cleaning and annotation, the R-RAD dataset includes a total of 3515 medical questions and 314 medical images, while the R-SLAKE dataset comprises 5980 medical questions and 546 medical images. Both the R-RAD and R-SLAKE datasets contain open-ended and closed-ended questions, with relevant statistics detailed in Table LABEL:tab:_dataset.
3.2 Dataset Cleaning and Annotation
In light of advancements of MLLMs, we integrated GPT-4V [?] into our data cleaning and annotation process, aiming to streamline the workflows. This integration can not only expedite data processing but also unearth subtleties often missed in manual cleaning and annotation practices. As mentioned earlier, our data come from SLAKE and VQA-RAD, two medical datasets manually created by medical experts. Nevertheless, with the aid of GPT-4V, we still identified some erroneous samples requiring further modifications. For instance, we find inconsistencies within the same medical image: the original question of the dataset “Is/Are the right hemidiaphragm normal?” was answered with “No”, whereas another question “Is this image normal?” received a “Yes” response. This apparent contradiction prompted us to seek further expert medical review for such cases, ensuring the reliability of our dataset.
After data cleaning, we utilize GPT-4V for data annotation, specifically in generating medical decision-making rationales for the items within the VQA-RAD and SLAKE datasets, as shown in Figure LABEL:fig:_overivew (a). This involves furnishing GPT-4V with the datasets’ images, questions, and corresponding correct answers. The purpose of these rationales is to enhance the model’s reasoning capabilities. Note that the rationale texts shall not contain exact correct answers, otherwise, the model could directly infer answers without reasoning. Therefore, we develop a fixed prompt to guide the generation process of GPT-4V. Our method carefully ensures that the rationales are informative yet do not directly disclose the correct answers, maintaining the integrity of the model’s learning and reasoning mechanisms. To ensure the quality of rationales, the domain expert check their validity and applicability. Any rationale not meeting criteria will be regenerated by GPT-4V. If a rationale generated by GPT-4V remains below standard even after five attempts, the domain expert will personally create an acceptable version, adhering to predefined criteria.
3.3 Dataset Analysis
We segment the generated rationales into individual words. After excluding common stop words, we conduct the frequency analysis of the remaining words. The results are visually represented in a word cloud (Figure LABEL:fig:_wordcloud). This analysis reveals that the high-frequency words in the rationales not only pertain to medical terminology, like “brain”, “chest” and “lung”, but also describe the morphological structures in medical imagery, exemplified by terms like “located”, “transverse”, and “density”. This pattern suggests that our generated rationales encompass both relevant medical knowledge and detailed descriptions of medical images. Such comprehensive information can potentially enhance the AI model’s performance in MedVQA tasks. Furthermore, we examine the length distribution of the rationales (Figure LABEL:fig:_distribution). We observe that the length of rationales, typically ranging between 60 and 110 words, is consistent across different organ-related questions. This consistency suggests that our annotation process does not vary in detail or depth when addressing diseases of various organs, thereby maintaining a balanced approach.
3.4 Problem Formulation
Formulation. In this paper, We denote the medical dataset as $\mathcal{D}=\{(I_{m},T_{m},A_{m},R_{m})\}^{M}_{m=1}$ where $M$ is the number of data samples. And the goal of the MedVQA task is to develop a mapping function $f(\cdot)$ that can generate textual answers in response to the medical questions, represented as:
$\{A,R\}=f(I,T),$ (1)
Here, $I$ denotes the medical image sourced from modalities such as X-ray, CT, or MRI. $T$ represents the natural language question pertaining to the medical image $I$. The output of the model $f(\cdot)$, represented as $\{A,R\}$, comprises two components. $A$ is the predicted textual answer, directly addressing the query posed in $T$. $R$, termed as “medical decision-making rationale”, offers a detailed justification for the answer $A$, elucidating an interpretative insight into how the model processes $I$ and $T$.
Loss Function. Given the input $X=\{I,T\}$, the model $f$ is trained by maximizing the likelihood of accurately predicting the target output $Y$ = $\{A,R\}$. The training involves a loss function, primarily the negative log-likelihood of correctly predicting subsequent tokens in the sequence $Y$, accumulated over all time steps. This is mathematically formulated as:
$L=-\sum_{n=1}^{N}\log p(Y_{n}|X,Y^{1:n-1}),$ (2)
In this context, $N$ represents the total number of tokens in the target answer $Y$, and $p(Y_{n}|X,Y^{1:n-1})$ denotes the conditional probability of correctly predicting the $n$-th token in $Y$, given the input $X$ and all preceding tokens $Y^{1:n-1}$ in the sequence. This loss function significantly improves the model’s capability to accurately forecast each token in the target output, thereby enhancing its overall predictive performance.
Model Architecture. The model architecture comprises five components, as shown in Figure LABEL:fig:_overivew (b): TextualEncoder, VisualEncoder, Cross Attention Network, Gated Fusion Network, and TextualDecoder. Notably, the TextualEncoder, VisualEncoder and TextualDecoder are all based on the Transformer architecture, renowned for its powerful learning and representational capabilities.
The TextualEncoder vectorizes the input question $T$ into the textual feature space, represented as $F_{T}\in\mathbb{R}^{n\times d}$, while the VisualEncoder transforms the input medical image $I$ into vision features $F_{I}\in\mathbb{R}^{m\times d}$. This can be expressed as:
$F_{\text{T}}$ $=\text{TextualEncoder}(T),$ (3)
$F_{\text{I}}$ $=\text{VisualEncoder}(I),$ (4)
where $n$ denotes the length of the input language text, and $d$ indicates the hidden dimension, $m$ represents the number of image patches.
Upon acquiring the textual representation $F_{T}$ and visual representation $F_{I}$, our model employs the Cross-Attention Network to facilitate interaction between these two modalities. The Cross-Attention Network computes the attention-guided visual feature $H_{attn}^{I}\in\mathbb{R}^{n\times d}$, which captures the relevant visual features corresponding to the textual query through the following operation:
$H_{\text{attn}}^{\text{I}}$ $=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V,$ (5)
where $Q$, $K$, $V$ correspond to the query, key, and value, derived from $F_{T}$, $F_{I}$, $F_{I}$, respectively.
Subsequently, the Gated Fusion Mechanism is utilized to dynamically combine the textual representation $F_{T}$ and the attention-guided visual feature $H_{attn}^{I}$. It determines the fusion coefficient $\lambda$ through a sigmoid-activated linear combination of the two modalities:
$\lambda$ $=\text{Sigmoid}(W_{l}F_{Q}+W_{v}H_{\text{attn}}^{\text{I}}),$ (6)
The fused output $F_{\text{fuse}}\in\mathbb{R}^{n\times d}$ is then computed as a weighted sum of $F_{T}$ and $H_{\text{attn}}^{\text{I}}$, moderated by $\lambda$:
$F_{\text{fuse}}$ $=(1-\lambda)\cdot F_{\text{T}}+\lambda\cdot H_{\text{attn}}^{\text{I}},$ (7)
Here, $W_{l}$ and $W_{v}$ are the model parameters that are learned during training to optimize the fusion of information between textual and visual streams. Finally, the fused output $F_{\text{fuse}}$ is fed into the TextualDecoder to generate the output $\{A,R\}$:
$\{A,R\}$ $=\text{TextualDecoder}(F_{\text{fuse}}),$ (8)
Three Generation Strategies. To investigate the impact of the medical decision-making rationale on the model performance in MedVQA tasks, we present three different generation strategies. These strategies are designed to guide the model in generating various forms of outputs, corresponding to different orders of the medical decision-making rationale in the process of generation. The methods are categorized as “Explanation”, “Reasoning” and “Two-Stage Reasoning”, as shown in Figure LABEL:fig:_overivew (c).
In the “Explanation” method, the answer $A$ is generated first, followed by the medical decision-making rationale $R$. In contrast, the “Reasoning” method reverses this order, generating $R$ before $A$. The “Two-Stage Reasoning” method follows a phased strategy, where two independent models are trained in distinct stages. The first stage focuses on using the medical question $T$ and the medical image $I$ to generate the intermediate result $R$. In the second stage, a different model takes $R$, along with $T$ and $I$, to derive the final answer $A$.
4 Experiments
4.1 Setting
Table 1: Performance of Our Methods on Open-End Questions in the R-RAD and R-SLAKE Datasets.
{tblr}
cells = c, row1 = font=, hline1,8 = 1pt,solid, cell21 = r=3, cell51 = r=3, vline2-3 = 1-4,5-7, vline3 = 3-4,6-7, hline1-2,8 = -, Dataset & Method Rouge-1 Rouge-2 Rouge-L BLEU-1 BLEU-2 BLEU-3 BLEU-4
R-RAD Explanation 50.2 20.2 29.5 38.3 22.9 14.0 8.8
Reasoning 49.8 20.3 29.3 37.8 22.7 14.0 8.9
Two-Stage Reasoning 49.1 19.9 28.7 37.7 22.5 13.9 8.8
R-SLAKE Explanation 53.1 22.7 31.7 39.2 24.1 15.4 9.9
Reasoning 53.5 22.8 32.1 39.5 24.3 15.5 10.0
Two-Stage Reasoning 53.2 23.1 32.0 39.5 24.5 15.8 10.3
Table 2: Accuracy (%) Comparison of Methods on Closed-End Questions in the R-RAD and R-SLAKE Datasets. Gray and red backgrounds indicate results for established and our methods.
{tblr}
cells = c, colspec = ccc, row1 = font=, row2-8 = bg=gray!10, row9-12 = bg=red!10, hline1,13 = 1pt,solid, vline2-3 = solid, Methods R-RAD R-SLAKE
MFB [?] 74.3 75.0
SAN [?] 69.5 79.1
BAN [?] 72.1 79.1
MEVF+SAN [?] 73.9 78.4
MEVF+BAN [?] 77.2 79.8
MMBERT [?] 77.9 -
PubMedCLIP [?] 79.5 82.5
w/o R 79.0 82.5
Reasoning 73.9 80.8
Two-Stage Reasoning 80.5 79.1
Explanation 83.5 86.3
In our framework, the encoder and decoder from UnifiedQA [?] are integrated as TextualEncoder($\cdot$) and TextualDecoder($\cdot$), respectively. Additionally, DETR [?] is employed as VisualEncoder($\cdot$). This integration facilitates the construction of our experimental model. Our experimental procedure is systematically divided into distinct phases. We initially evaluate the performance of the “Explanation”, “Reasoning” and “Two-Stage Reasoning” methods on the R-RAD and R-SLAKE datasets, benchmarking them against established models like PubMedCLIP [?], MMBERT [?], MEVF [?], BAN [?], SAN [?], MFB [?]. Subsequently, the quality of the rationales generated through our proposed methods, was assessed.
In our experiments, the learning rate is uniformly set at 5e-4 for both the R-SLAKE and R-RAD datasets. However, the number of epochs in the fine-tuning phase differs, comprising 300 epochs for the R-SLAKE dataset, but only 150 epochs for the R-RAD dataset. It is important to note that our “Two-Stage Reasoning” method requires a phased fine-tuning process involving two separate models. For the first model, we follow the parameters mentioned earlier. For the second model, we consistently use a learning rate of 5e-5, fine-tuning it for 20 epochs for both datasets. All experiments reported in this paper are conducted using PyTorch on an Ubuntu server equipped with eight NVIDIA RTX 3090 GPUs. The code and datasets from our study will be released.
4.2 Main Results
Our performance evaluation is divided into two parts, focusing separately on closed-end and open-end questions. Closed-end questions, structured as multiple-choice questions with a single definitive answer, are assessed using accuracy as the performance metric, as shown in Table 4.1. In contrast, open-end questions allow for a range of answers due to their inherent nature. The answers generated by our methods are difficult to match precisely against the dataset. Therefore, we employ text generation metrics like Rouge and BLEU to evaluate our methods’ performance, as exhibited in Table 4.1.
In facing closed-end questions, the “Explanation” method surpasses the “Reasoning” and “Two-Stage Reasoning” methods with accuracy of 83.5% on the R-RAD dataset and 86.3% on the R-SLAKE dataset, marking improvements of 4.0% and 3.8%, respectively, over the state-of-the-art PubMedCLIP model.
With open-end questions, different methods show distinct advantages. The Rouge scores, similar to the “Recall”, emphasizes the completeness of the generated text, while the BLEU scores, akin to the “Precision”, stresses the preciseness of the generated text. The “Explanation” method demonstrates higher Rouge and BLEU scores on the R-RAD dataset, with Rouge-1, Rouge-L, BLEU-1, BLEU-2, and BLEU-3 reaching 50.2%, 29.5%, 38.3%, 22.9%, and 14.0%, respectively. The “Two-Stage Reasoning” method showcases higher scores on the R-SLAKE dataset, with Rouge-2, BLEU-1, BLEU-2, BLEU-3, and BLEU-4 at 23.1%, 23.1%, 39.5%, 24.5%, 15.8%, and 10.3%, respectively. The “Reasoning” method maintains robust performance across both the R-RAD and R-SLAKE datasets; on the R-RAD dataset, Rouge-2, BLEU-3, and BLEU-4 reached 20.3%, 14.0%, and 8.9%, respectively, while on the R-SLAKE dataset, Rouge-1, Rouge-L, and BLEU-1 are 53.5%, 32.1%, and 39.5%, respectively. These outcomes highlight the necessity of diverse methods for various types of open-end questions.
4.3 Rationale Quality Assessment
To further compare the “Explanation”, “Reasoning”, and “Two-Stage Reasoning” methods, we design two sets of experiments for closed-end questions on the R-RAD and R-SLAKE datasets. In the first set, we introduce a control experiment “w/o R”, where the models are trained and inferred without incorporating medical decision-making rationales. The “Explanation”, “Reasoning” and “Two-Stage Reasoning” methods are compared with the control experiment. As indicated in Table 4.1, relative to “w/o R”, the “Explanation”, “Reasoning”, and “Two-Stage Reasoning” methods showed accuracy improvements of 4.5%, -5.1%, and 1.5% respectively on the R-RAD dataset, and 3.8%, -1.7%, and -3.4% respectively on the R-SLAKE dataset.
In the second experimental series, we employ Gemini Pro to evaluate the assistance of medical decision-making rationales generated through our methods. Initially, the input of Gemini Pro is restricted to medical queries and associated imagery. Subsequently, medical decision-making rationales generated by the “Explanation”, “Reasoning” and “Two-Stage Reasoning” methods are incorporated to assist Gemini Pro in answering. We apply the self-consistency protocol [?], where Gemini Pro answers each question five times, selecting the most frequent answer as the final answer. The results, presented in Table LABEL:tab:_gemini_closed-end_acc, indicate an initial accuracy of 73.2% on the R-RAD dataset and 72.8% on the R-SLAKE dataset for Gemini Pro. The integration of medical decision-making rationales has led to significant improvements. Among three methods, the “Explanation” method stands out, enhancing the accuracy by 8.8% on the R-RAD dataset and 8.5% on the R-SLAKE dataset.
4.4 Ablation Study
To validate the impact of MLLM selection and expert annotation involvement on data quality during the data annotation process, we design three variations for annotating the closed-end questions in the R-RAD dataset: using Gemini Pro without expert input, using GPT-4V without expert input, and using GPT-4V with expert input. The results, as depicted in Figure LABEL:fig:_expert, indicate that GPT-4V enhances MedVQA task performance over Gemini Pro, aligning with prior researches [??] that illustrate GPT-4V’s more advanced reasoning abilities. Additionally, the participation of expert annotations is found to improve data quality.
4.5 Case Study
To observe the assistance of medical decision-making rationales in MedVQA tasks specifically, Figure LABEL:fig:_case_study shows several examples where the model employs the “Explanation” method. For instance, when addressing the question “Are the small bubbles of air seen abnormal?”, the model without Medical Decision-Making Rationales incorrectly responds. In contrast, the model utilizing the “Explanation” method not only provides the correct answer but also supplies the requisite medical knowledge, stating “The small bubbles appear as very dark or black on the CT scan”. Furthermore, it delivers pertinent descriptions of the medical image, noting “there are multiple areas of radiolucency, which are consistent with the appearance of small bubbles of air”. This underscores the supportive role of medical decision-making rationales in guiding the model to answer MedVQA questions accurately.
5 Conclusion
In this paper, we present a generative model-based framework for MedVQA and construct the R-RAD and R-SLAKE datasets, which include intermediate reasoning steps to address the challenge of black-box decision-making processes in MedVQA models. Extensive experimental results show that our proposed framework not only elucidates the medical decision-making process of MedVQA models with clarity but also significantly enhances their performance. Future research will further explore generative models tailored for real clinical settings and how to better evaluate the performance of MedVQA models in open-ended scenarios.